The perimeter of a quarter circle is 17.85 what is the circles radius
1
Expert's answer
2020-04-06T16:23:50-0400


P=(2"\pi"R)/4+2R is the perimeter of a quarter circle. R is the radius of the circle,

P=πR/2+2R=R(π/2+2)=17.85,

R(π/2+2)=17.85,

R=17.85/(π/2+2)=17.85/(3.14/2+2)=5.


So, the radius of the circle is R=5.